9) Six litres of 0.4M 3-(N-morpholino) propane sulfonic acid (MOPS) buffer of pH 7.8 are required. Available in the laboratory are 10M HCl and 10M NaOH and the basic form of MOPS (M.Wt = 210.3). The pKa for the MOPS zwitter ion is 7.2. Describe, showing the relevant calculations, the procedure for preparing the buffer
1
Expert's answer
2019-09-11T05:23:20-0400

The pH of this buffer should be calculated with the following expression:


"pH=pK_{a}+lg(\frac{C_{BA}}{C_{HA}})", MOPS is a middle strength acid (pKa = 7.2).


We can easily extract the required relation from the formula:


"\frac{C_{BA}}{C_{HA}}=10^{pH-pK_{a}}=10^{7.8-7.2}=3.98"


On the other hand, the required concentration of the buffer is 0.4M:


"C_{BA}+C_{HA}=0.4"


Now, we can calculate the required concentrations of MOPS and its Na-salt:


"3.98*C_{HA}+C_{HA}=0.4"


"C_{HA}=\frac{0.4}{3.98+1}=0.08M"


"C_{BA}=0.4-C_{HA}=0.32M"


"n(HA)=C_{HA}*V_{sol}=0.08M*6L=0.48mol"


"n(BA)=C_{BA}*V_{sol}=0.32M*6L=1.92mol"


"n(HA)+n(BA)=0.48mol+1.92mol=2.40mol"


You have to take 2.40mol*209.3g/mol = 502g of MOPS and 1.92mol/10M = 0.192L of 10M solution of NaOH (react in a 1 to 1 molar relation):


"C_{7}H_{15}NO_{4}S+NaOH \rightarrow C_{7}H_{14}NO_{4}SNa+H_{2}O"


When it is complete, you have to dissolve the mixture up to 6L in a special analytical flask with using of double distilled water.